# An Economical Approach to Design with Precision Criteria

### v3

@@ 7 Discussion
Our proposed framework can be broadly applied to design studies with precision criteria when CIs are constructed using a method that is based on likelihood functions. Future work could extend this methodology to accommodate design with CIs that are based on the frequentist nonparametric or the Bayesian bootstrap. Our approach could also be extended in other aspects to accommodate more complex design criteria. For instance, it would be useful to consider economical sample size determination methods for studies that aim to satisfy criteria for both the study power and the precision of interval estimates.
Additionally, we could further develop design methods with precision criteria in a sequential framework or in a framework with multiple targets of inference. Such extensions would require additional theoretical development because we would need to account for the level of dependence in the joint sampling distribution of CI lengths across all potential analyses and targets of inference. Finally, our design framework requires that the regularity conditions for the asymptotic normality of the MLE and BvM theorem are satisfied. It would therefore be of interest to consider design methods with precision criteria in settings where these regularity conditions are relaxed.

### v4

@@ 7 Discussion
Our proposed framework can be broadly applied to design studies with precision criteria when CIs are constructed using a method that is based on likelihood functions. Nevertheless, similar theory and methods can be developed for certain types of interval estimates that are not likelihood based. For instance, suppose the target of inference $\theta$ is a function of the regression parameters $\boldsymbol{\eta}$ from a generalized estimating equation (GEE). The variance of the sampling distribution of the M-estimator $\hat{\theta}^{{}_{(n)}}_{r}$ takes the form $n^{-1}\Lambda(\boldsymbol{\eta}_{r})$ (Liang and Zeger, 1986), where $n$ is the number of ISUs and $\Lambda(\boldsymbol{\eta}_{r})$ is based on an ISU-level sandwich variance estimator instead of $\mathcal{I}_{\emph{ISU}}(\boldsymbol{\eta}\mid{\boldsymbol{x}})$. A similar result to Theorem 1 can therefore be developed for use with GEEs such that $\mu_{l}$ and $\sigma_{l}$ are functions of the sandwich variance instead of the Fisher information. Because Algorithm 2 requires empirical estimation of CI lengths, it can be implemented as is with interval estimates for GEE regression parameters.
Future work could extend this methodology to accommodate semiparametric and nonparametric interval estimates more broadly, including design with CIs that are based on the frequentist nonparametric or the Bayesian bootstrap. It would be important to formally define settings under which analogues to Theorem 1 are theoretically valid to substantiate the use of our methods. While semiparametric and nonparametric methods can allow for more flexibility in data analysis, we emphasize that we still need to fully specify data generation processes to implement simulation-based design in these settings. Our approach could also be extended in other aspects to accommodate more complex design criteria. For example, it would be useful to consider economical sample size determination methods for studies that aim to satisfy criteria for both the study power and the precision of interval estimates.